OpenStudy (anonymous):

Choose the correct product of (2x − 5)2. does the power change the sign ? like would it be 4x^2-10 or 4x^2+25

OpenStudy (foolaroundmath):

Neither. Try multiplying term by term for this (2x-5)(2x-5) = 2x(2x-5) -5(2x-5)

OpenStudy (anonymous):

4x^2-20x+25 then?

OpenStudy (anonymous):

Here you have to use the formula: \[\huge \color{green}{(a-b)^2 = a^2 + b^2 - 2ab}\]

OpenStudy (anonymous):

Yes, This one is right...

OpenStudy (foolaroundmath):

Yes, the answer is right. I was trying to point towards the identity that waterineyes posted. It's something that should be remembered. (its easy enough to derive using the same procedure of multiplying term by term) \((a-b)^{2} = (a-b)(a-b)\)